The mass of the planet Mars is 6.39 × 1023 kg. Its radius is 3.39 × 106 m. The gravitational field strength on the surface of Mars is:

(a) 0.27 N kg–1

(b) 3.71 N m–2

(c) 0.371 N kg–1

(d) 3.71 N kg–1


(1 Mark)


1
Expert's answer
2021-02-19T10:30:19-0500
"g=\dfrac{GM}{r^2}=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot6.39\cdot10^{23}\ kg}{(3.39\cdot10^6\ m)^2}=3.71\ \dfrac{N}{kg}."

Hence, the correct answer is (d).

Answer:

(d) 3.71 N kg–1